The Lawrence Sisters Moving Company paid a quarterly dividend of $2.06 per share last quarter. Today, the company announced that future dividends will be increasing by 1.4% percent quarterly. If you require a 10.07% percent annual rate of return, how much are you willing to pay to purchase one share of this stock today? (Round intermediate steps to 4 digits.)

Select one:

a. $ 180.75

b. $ 192.39

c. $ 192.55

d. $ 186.92

e. $ 180.43
Answer #1

Dividend = $2.06

Growth rate = 1.4%

Quarterly rate of return = 0.1007 / 4 = 0.025175 = 2.5175%

Price of stock = Dividend * (1 + Growth rate) / (Rate of return - Growth Rate)

Price of stock = 2.06 * (1.014) / (0.025175 - 0.014)

Price of stock = 186.9208 = $186.92